# Faithful, Interpretable Chest X-ray Diagnosis with Anti-Aliased B-cos Networks

### v1

@@ Abstract
Faithfulness and interpretability are essential for deploying deep neural networks (DNNs) in safety-critical domains such as medical imaging. B-cos networks offer a promising solution by replacing standard linear layers with a weight-input alignment mechanism, producing inherently interpretable, class-specific explanations without post-hoc methods. While maintaining diagnostic performance competitive with state-of-the-art DNNs, standard B-cos models suffer from severe aliasing artifacts in their explanation maps, making them unsuitable for clinical use where clarity is essential. Additionally, the original B-cos formulation is limited to multi-class settings, whereas chest X-ray analysis often requires multi-label classification due to co-occurring abnormalities. In this work, we address both limitations: (1) we introduce anti-aliasing strategies using FLCPooling (FLC) and BlurPool (BP) to significantly improve explanation quality, and (2) we extend B-cos networks to support multi-label classification. Our experiments on chest X-ray datasets demonstrate that the modified and preserve strong predictive performance while providing faithful and artifact-free explanations suitable for clinical application in multi-label settings. Code available at: GitHub repository.
@@ 1 Introduction
Faithfulness and interpretability are critical prerequisites for deploying deep learning models in safety-critical domains such as healthcare, where decisions have direct implications for patient outcomes [49, 61]. In particular, clinical adoption demands models whose reasoning processes can be understood and verified by medical professionals. However, most existing approaches in medical image processing rely on architectures that lack inherent interpretability, offering limited insight into the basis of their predictions [42]. As emphasized in [61], knowing how a diagnosis was derived is highly valuable, especially in practice, where AI tools are used to support, not replace, radiologists. In this work, we explore B-cos networks [8], which are inherently interpretable by design and generate class-specific contribution maps that visually indicate the evidence behind each prediction. Representative examples are shown in Fig. 1 which compares our artifact-free, faithful and interpretable explanations to other common approaches such as GradCAM and LayerCAM while displaying the need for our extensions when working with B-cos networks and chest X-rays simultaneously.
These models can directly show which part of the image caused network activations, enabling an understanding of why a classification was made [8]. Unlike post-hoc approaches such as GradCAM [50] or LayerCAM [29], which provide coarse and sometimes misleading heatmaps that have insufficient mechanistic explanation of the decision-making process [17, 48], B-cos networks offer inherently interpretable, class-specific contribution maps that yield clearer, more faithful visual explanations. This can help radiologists verify predictions, improve clinical workflows, and reduce diagnostic errors. However, standard B-cos models face two critical limitations: first, the generated explanation maps often exhibit grid artifacts, making them visually unreliable for clinical use; second, the original formulation supports only multi-class classification, whereas chest X-rays commonly involve multiple co-occurring conditions. In this work, we address both limitations by incorporating anti-aliasing techniques such as FLCPooling and BlurPool to improve visual clarity, and by extending the B-cos framework to support multi-label classification. We refer to the resulting anti-aliased, multi-label capable model as and . While combines B-cos models with FLCPooling [22], combines it with BlurPooling [59], both effective anti-aliasing techniques. These anti-aliasing techniques help improve the explanation maps from the B-cos network, since they replace the artifact-producing downsampling operation with an artifact-free downsampling path. Thus, removing the artifacts in the explanation maps of B-Cos networks. Our main contributions are:
@@
An extension of the B-cos framework to multi-label classification, addressing the clinical need to detect multiple co-existing conditions in a single chest radiograph.
@@ 2 Related Work
Model Architectures: Convolutional neural networks (CNNs) remain the most widely used architectures for chest X-ray classification tasks. ResNet [24], and specifically ResNet50 [24], has been effectively applied to COVID-19 detection on large-scale datasets such as CheXpert [28] and NIH ChestX-ray14 [57], showing high performance [25]. CheXNet [45], based on DenseNet-121 [27], reports a pneumonia detection F1-score of 0.435, surpassing the average radiologist performance of 0.387 on the same dataset [57]. Lightweight CNN variants such as MobileNet [26] and EfficientNet [46] have also demonstrated competitive results across several datasets [47, 46]. More recently, vision transformers (ViTs) [53]. IEViT [43], a ViT variant designed specifically for medical imaging, further improves upon standard ViTs and achieves F1-scores of up to 100% across four chest X-ray datasets [43].
@@
For this work, we focus on datasets that provide region-level annotations, which are essential for quantitative assessment of visual explanations. In the single-label setting, we use the RSNA Pneumonia Detection Challenge (2018) dataset [51], which contains 26,684 radiographs annotated as either pneumonia or no pneumonia, along with bounding boxes localizing the affected regions. Despite being designed for a Kaggle challenge, the dataset remains widely used for both image classification and object detection tasks due to its high-quality annotations. For the multi-label setting, we use the VinBigData Chest X-ray Abnormalities Detection dataset, a modified version of VinDr-CXR [41], consisting of 15,000 radiographs with bounding boxes for 14 thoracic abnormalities. This dataset is particularly valuable because all annotations were manually curated by 17 board-certified radiologists [41], ensuring high-quality supervision. Moreover, it reflects realistic clinical cases where multiple pathologies often co-occur within a single image, making it well-suited for the multi-label classification task studied in this work.
@@ 3.1 Extending B-cos for Multi-label Classification
B-cos networks [8] provide inherently interpretable outputs by computing class-specific contribution maps, but are originally designed for multi-class classification. In such cases, only the explanation map corresponding to the predicted class is generated. However, in medical imaging, multiple pathologies often co-occur, making multi-label classification essential. To address this, we extend B-cos to compute contribution maps for all output neurons simultaneously. This enables a full analysis of model behavior across all conditions and does not alter the underlying training or inference pipeline. Instead, it enhances the interpretability by allowing inspection of every class output, which is critical in the medical domain where several diseases can co-occur.
@@ 4.2.1 Interpretability Evaluation
Turning to the multi-label setting, results on the VinBigData dataset are shown in Tab. 6. Our best-performing models outperform the established CheXNet [45], which achieved an F1-score of 0.435 on a comparable 14-class task. These results validate the applicability of our B-cos-based models beyond binary classification.
@@ 4.3 Qualitative Results
In addition to evaluating raw B-cos explanations, it is instructive to compare them with post-hoc methods such as GradCAM. As shown in Fig. 7, B-cos maps capture finer structural detail and more precisely delineate the decision-relevant regions compared to GradCAM heatmaps, which often appear diffuse and ambiguous. The red bounding boxes in the unhealthy example also confirms strong alignment with disease indicators from our model. The spatial precision of B-cos outputs can aid radiologists in verifying predictions and identifying subtle indicators of disease, making them more suited for medical decision support [8].
@@ 6 Conclusion
This work provides a comprehensive analysis of B-cos networks for chest X-ray classification, targeting two key limitations that have restricted their application in medical imaging: aliasing artifacts in explanation maps and the lack of support for multi-label classification. We address the former by integrating anti-aliasing techniques, FLCPooling, and BlurPool, which improve the clarity and reliability of the explanations. We address the latter by extending B-cos to produce class-wise contribution maps in multi-label settings, making the method suitable for realistic diagnostic tasks. Our experiments on the RSNA Pneumonia Detection Challenge and VinBigData Chest X-ray Abnormalities Detection datasets confirm that B-cos networks provide competitive classification performance while offering inherently interpretable, high-fidelity explanations. These explanations align closely with disease-relevant regions, as demonstrated both qualitatively and through strong energy-based pointing game scores, outperforming LayerCAM and established models such as CheXNet. By enhancing the interpretability and practical applicability of B-cos networks, this work takes an important step toward transparent and trustworthy AI-assisted diagnosis in clinical settings [61], and sets the stage for future research on scalable, interpretable deep learning for medical imaging.

### v2

@@ Abstract
Faithfulness and interpretability are essential for deploying deep neural networks (DNNs) in safety-critical domains such as medical imaging. B-cos networks offer a promising solution by replacing standard linear layers with a weight-input alignment mechanism, producing inherently interpretable, class-specific explanations without post-hoc methods. While maintaining diagnostic performance competitive with state-of-the-art DNNs, standard B-cos models suffer from severe aliasing artifacts in their explanation maps, making them unsuitable for clinical use where clarity is essential. In this work, we address these limitations by introducing anti-aliasing strategies using FLCPooling (FLC) and BlurPool (BP) to significantly improve explanation quality. Our experiments on chest X-ray datasets demonstrate that the modified and preserve strong predictive performance while providing faithful and artifact-free explanations suitable for clinical application in multi-class and multi-label settings. Code available at: GitHub repository.
@@ 1 Introduction
Faithfulness and interpretability are critical prerequisites for deploying deep learning models in safety-critical domains such as healthcare, where decisions have direct implications for patient outcomes [50, 64]. In particular, clinical adoption demands models whose reasoning processes can be understood and verified by medical professionals. However, most existing approaches in medical image processing rely on architectures that lack inherent interpretability, offering limited insight into the basis of their predictions [42]. As emphasized in [64], knowing how a diagnosis was derived is highly valuable, especially in practice, where AI tools are used to support, not replace, radiologists. In this work, we explore B-cos networks [8], which are inherently interpretable by design and generate class-specific contribution maps that visually indicate the evidence behind each prediction. Representative examples are shown in Fig. 1, which compares our artifact-free, faithful, and interpretable explanations to other common approaches such as GradCAM and LayerCAM while displaying the need for our extensions when working with B-cos networks and chest X-rays simultaneously.
These models can directly show which part of the image caused network activations, enabling an understanding of why a classification was made [8]. Unlike post-hoc approaches such as GradCAM [51] or LayerCAM [29], which provide coarse and sometimes misleading heatmaps that have insufficient mechanistic explanation of the decision-making process [17, 49], B-cos networks offer inherently interpretable, class-specific contribution maps that yield clearer, more faithful visual explanations. This can help radiologists verify predictions, improve clinical workflows, and reduce diagnostic errors [56, 57]. However, standard B-cos models face a critical limitation: the generated explanation maps often exhibit grid artifacts, making them visually unreliable for clinical use. In this work, we address these limitations by incorporating anti-aliasing techniques such as FLCPooling and BlurPool to improve visual clarity. Additionally, while the original B-cos models support multi-label classification, they were proposed for multi-class classification. Thus, we use the framework from [46], to obtain multi-label classification since chest X-rays commonly involve multiple co-occurring conditions. We refer to the resulting anti-aliased, multi-label capable model as and . While combines B-cos models with FLCPooling [22], combines it with BlurPooling [62], both effective anti-aliasing techniques. These anti-aliasing techniques help improve the explanation maps from the B-cos network, since they replace the artifact-producing downsampling operation with an artifact-free downsampling path. Thus, removing the artifacts in the explanation maps of B-Cos networks. Our main contributions are:
@@
We show that our proposed framework can be adopted for both multi-class and multi-label classification, proving useful in critical medical applications.
@@ 2 Related Work
Model Architectures: Convolutional neural networks (CNNs) remain the most widely used architectures for chest X-ray classification tasks. ResNet [24], and specifically ResNet50 [24], has been effectively applied to COVID-19 detection on large-scale datasets such as CheXpert [28] and NIH ChestX-ray14 [60], showing high performance [25]. CheXNet [45], based on DenseNet-121 [27], reports a pneumonia detection F1-score of 0.435, surpassing the average radiologist performance of 0.387 on the same dataset [60]. Lightweight CNN variants such as MobileNet [26] and EfficientNet [47] have also demonstrated competitive results across several datasets [48, 47]. More recently, vision transformers (ViTs) [54]. IEViT [43], a ViT variant designed specifically for medical imaging, further improves upon standard ViTs and achieves F1-scores of up to 100% across four chest X-ray datasets [43]. Lastly, [57] uses B-cos models for interpretable and explainable predictions for cellular data from the medical domain, consolidating the validity of our approach.
@@
For this work, we focus on datasets that provide region-level annotations, which are essential for the quantitative assessment of visual explanations. In the single-label setting, we use the RSNA Pneumonia Detection Challenge (2018) dataset [52], which contains 26,684 radiographs annotated as either pneumonia or no pneumonia, along with bounding boxes localizing the affected regions. Despite being designed for a Kaggle challenge, the dataset remains widely used for both image classification and object detection tasks due to its high-quality annotations. For the multi-label setting, we use the VinBigData Chest X-ray Abnormalities Detection dataset, a modified version of VinDr-CXR [41], consisting of 15,000 radiographs with bounding boxes for 14 thoracic abnormalities. This dataset is particularly valuable because all annotations were manually curated by 17 board-certified radiologists [41], ensuring high-quality supervision. Moreover, it reflects realistic clinical cases where multiple pathologies often co-occur within a single image, making it well-suited for the multi-label classification task studied in this work.
@@ 3.1 Extending B-cos for Multi-label Classification
B-cos networks [8] provide inherently interpretable outputs by computing class-specific contribution maps. While they are capable of multi-label classification as teased by [8], they are originally designed for multi-class classification. In such cases, only the explanation map corresponding to the predicted class is generated. However, in medical imaging, multiple pathologies often co-occur, making multi-label classification essential. To address this, we adopt the B-cos multi-label models from [46] to compute contribution maps for all output neurons simultaneously. This enables a full analysis of model behavior across all conditions and does not alter the underlying training or inference pipeline. Instead, it enhances the interpretability by allowing inspection of every class output, which is critical in the medical domain where several diseases can co-occur.
@@ 4.2.1 Interpretability Evaluation
Next, the multi-label setting results on the VinBigData dataset are shown in Tab. 6. Our best-performing models outperform the established CheXNet [45], which achieved an F1-score of 0.435 on a comparable 14-class task. These results validate the applicability of our B-cos-based models beyond binary classification.
@@ 4.3 Qualitative Results
In addition to evaluating raw B-cos explanations, it is instructive to compare them with post-hoc methods such as GradCAM. As shown in Fig. 7, B-cos maps capture finer structural detail and more precisely delineate the decision-relevant regions compared to GradCAM heatmaps, which often appear diffuse and ambiguous. The red bounding boxes in the unhealthy example also confirm strong alignment with disease indicators from our model. The spatial precision of B-cos outputs can aid radiologists in verifying predictions and identifying subtle indicators of disease, making them more suited for medical decision support [8].
@@ 6 Conclusion
This work provides a comprehensive analysis of B-cos networks for chest X-ray classification, targeting key limitations that have restricted their application in medical imaging: aliasing artifacts in explanation maps. We address them by integrating anti-aliasing techniques, FLCPooling, and BlurPool, which improve the clarity and reliability of the explanations. We demonstrate that the anti-aliasing methods work with both multi-class and multi-label B-cos models to produce contribution maps, making the method suitable for realistic diagnostic tasks. Our experiments on the RSNA Pneumonia Detection Challenge and VinBigData Chest X-ray Abnormalities Detection datasets confirm that B-cos networks provide competitive classification performance while offering inherently interpretable, high-fidelity explanations. These explanations align closely with disease-relevant regions, as demonstrated both qualitatively and through strong energy-based pointing game scores, outperforming LayerCAM and established models such as CheXNet. By enhancing the interpretability and practical applicability of B-cos networks, this work takes an important step toward transparent and trustworthy AI-assisted diagnosis in clinical settings [64], and sets the stage for future research on scalable, interpretable deep learning for medical imaging.